OLIS Logbook General
  General, Page 1 of 26  Not logged in ELOG logo
Entry  Wed Feb 22 21:51:40 2017, Keerthi Jayamanna, , , SIS first beam is N1+1 

Chris confirmed the first SIS beam would be Na+1. Next Rb+1 and then Li+1

Entry  Wed Feb 22 21:46:27 2017, Keerthi Jayamanna, , , Elog is working now 

Elog is now functioning. We can start entering notes

Entry  Mon Sep 14 11:21:16 2015, Keerthi Jayamanna, Routine, General, 20Ne+2 beam saved 

20Ne+2 beam optics saved as 150914_1117_ios0iosfc6.snap

Entry  Thu May 21 09:15:48 2015, Brian Minato, , ,  

Service of MCIS.  Plasma chamber cleaned.  Repaired shorted, Right steering plate.  Added shields to help prevent coating of wires.  Beam is off centre on first collimator of steering assembly.

Moved Ground electrode away from PE approximately 0.10".

  Note:  the plasma chamber and ground electrode tip are clean.

2015-05-19  Repaired shorted MWS PE.  Internal short was found.  The source chamber was vented and cleaned.  The beaker was replaced with a new one (it was coated with carbon).  WPI2015-05-19-2

Investigate Shorted MWS PE and MCIS shorted steering plate WOI2015-05-19-1

Entry  Wed Apr 29 07:48:22 2015, Brian Minato, Other, Other, Repair- Water leak inside HV cage 

A water leak was reported inside the HV cage (FR#8079).  The cage was accessed to determine the source of the leak (WP#I2015-04-27-1).

The WMS cusp cooling line was leaking.  It was close to the MCIS cart (at ground) and had sparked through.  The hose was cut at the puncture location and re-connected. to the fitting.  It was re-routed to be further away from the MCIS cart.

Entry  Mon Apr 20 08:09:03 2015, Brian Minato, Routine, General, OLIS Shutdown Update 

The OLIS 1 week shutdown has been completed (WP 2015-04-13-1).


Y-Box insulators were inspected and look good.  Beamline optics were inspected.


The La disk was removed.  Source housing was cleaned and re-asssembled with a beaker installed. 


The cart was installed in the HV cage.  Kr gas pack was installed on the cart and the transfer tube was opened up to MCIS vacuum.


Source PE and EE assembly was removed and opened up for inspection.  A new source was assembled (1/4 Ta tube and 0.040 Ta Oven & Ionizer wires) and installed.  Lithium Nitrate Anhydrous (ALFA #13405 TR#205938) was inserted in the oven.


Additional details-



OLIS MWS is currently running an producing 22Ne2+


Entry  Wed Dec 17 13:17:13 2014, Brian Minato, Other, Other, MCIS Update 

The MCIS was vented with nitrogen yesterday.  The Se oven was drawn back into the tuner.  It was removed with the tuner and blanked.

The standard quartz window was removed and replaced with a ceramic, test window supplied by CPI. (p/n VWU2054).  The system is currently pumping down.

Entry  Tue Nov 25 12:23:28 2014, Keerthi Jayamanna, Routine, General, 76Se+13 tune saved 

141125_1159_OLIS_76Se+13.snap is saved in reference directory.





Frequency = 14.44765GHz

Amplitude = -10.880

TET FP= 9.8W

TWT RP = 1.8W


Entry  Tue Nov 25 07:49:50 2014, Brian Minato, Other, General, MCIS and SIS Update 


the Se was loaded into the oven last friday (21st).  During the weekend the presence of 76Se +11 to +15 was verified.

NOTE- the oven p/s on the cart is broken.  An attempt was made to remove it on friday.  It may be possible to remove, if the rack it is mounted on is loosened and turned.



the Ionizer p/s control cable was removed to allow for local operation.  Keerthi tested the power supply in local mode on the weekend and found that the voltage did fluctuate.  This was confirmed by changes in the beam current.

Set-up pictures are located in -

M:\OLIS\OFF LINE ION SOURCE\SURFACE SOURCE\Single Source - Purchased Box\PICTURES\2014-11-21 Voltage Fluctuation setup

Ionizer Power supply + is connected to the common on the SIS source

Ionizer Power supply - is connected to the Ionizer on the SIS source

a cable is connected from the SIS main, bias plate to the negative terminal on the p/s.

Entry  Wed Nov 19 11:34:55 2014, Raymond Dube, Routine, General, OLIS Electrical diagram with 3 I.S. ISK4586D-rev2.pdf

I updated the electrical Diagram  for OLIS Faraday Cage with 3 Ion Source, ref: ISK4586D rev. 2

PS:  Make sure that you do not connect both Bias P/S on same Ion Source ...

Entry  Tue Nov 18 08:56:23 2014, Brian Minato, Other, General, ALGA Update 

There has been no noticeable loss of communication to the ALGA 50watt amplifier.  As a test, the ethernet cable was removed for about a minute and then reconnected.  The amplifer re-established communication without issue.  This test was done twice.

Now, the CPI ethernet cable has been returned to the CPI amplifier.  The original ALGA ethernet cable (blue) has been reconnected to the ALGA.

PS  The nitrogen bottle has been replaced.

Entry  Thu Nov 13 16:34:29 2014, Brian Minato, Other, Other, MCIS Update 

The ethernet cable from the CPI TWT amplifier has been temporarily moved to the Alga 50W amplifier.  The Alga had been experiencing loss of communication.  This has not been an issue with the CPI so this cable will be tried with the Alga.  A signal generator is not connected to the Alga.

Entry  Thu Nov 13 16:31:54 2014, Brian Minato, Other, Other, SIS 

A new SIS source has been installed.  The oven has been loaded with Sodium Hydrogen Carbonate (Alfa Aesar 10863) and Rubidium Nitrate.

Pictures are located in 2014-11-13 Source and PS Pics

Entry  Wed Nov 12 13:32:47 2014, Raymond Dube, Routine, General, New in-line resistors for IOS:PE ISK4583D-rev10.pdf

Today we replace the in-line resistors for the Plasma Electrode (PE) it was a 25 ohm 200W (2 x 50 Ohm 100W) with a 50 Ohm 200W (2 x 100 Ohm / 100W) also added the inductance coil.

This is R5 on drawing: ISK4583D rev. 10

This was done to reduce the power on R5

Note: We still need to make sure the the FAN is running to cool the resistors when the PE P/S is ON.

Entry  Thu Nov 6 08:37:52 2014, Brian Minato, Other, Other, OLIS Update 


the source is off.  It is under vacuum and in the HV cage.


the test filament assembly has been removed for inspection and the port blanked.  Ray D. noticed that the Swagelok connections for the water manifold are extremely close to each other (potential short).  Yesterday, when reducing the current on the ionizer the voltage varied significantly.  The currently on the FC also changed.

The temporary collimator installed in the 4-1/2 CF port has been modified.  the through hole opening has been increased from 3mm diameter to 6mm diameter.


Entry  Sat Nov 1 13:31:17 2014, Keerthi Jayamanna, , , MCIS readyness for 76Se 

Supernanogan is running with oven no heat, without 76selenium. MCIS:IG1=3e-7Torr.

Tune saved as 141101_1318_76Se+11tune.snap

MCIS steering  X=330, Y=226, CCB1=500, MCISEL=4171/5157, PE=11.5/23, FC6=76nA(16O+7)

Entry  Fri Oct 31 20:39:56 2014, Brian Minato, Other, Other, OLIS Update 


the MWS Bias and PE have been moved on to the MCIS (the MCIS connections have been removed).

The plug in power to the HP Signal generator has been moved to a continuous source (Removed from the source controlled by the SSW).


a test filament assembly has been installed in the SIS.  No chemicals have been loaded or alignment done on the oven / ionizer.  It has been installed just as a test.

Entry  Thu Oct 30 23:44:57 2014, Keerthi Jayamanna, Routine, General, OLIS Supernanogan mass scan 6 to 7.5 

A/q           FC6  (only current over 100pA)

6.0           1.85e-7

6.15         4.96e-10

6.23         2.4e-10

6.31         6.14e-10

6.39         2e-10

6.51         1.82e-9

6.90         2.7e-10

7.00         5e-8

7.1           4.6e-9

7.27        1e-10

7.5        3.32e-10


Entry  Thu Oct 30 23:37:58 2014, Keerthi Jayamanna, Routine, General, Mass scan 4 to 6 

A/Q            FC6

4                1.1e-6

4.25          8e-10

4.34          1.2e-9

4.42          5.5e-9

4.51          2.59e-9

4.58          1.7e-10

4.66          1.7e-10

4.74          1.7e-10

4.85          4.6e-10

5.0            1.36e-10

5.07          1.5e-10

5.13         2.69e-10

5.25         4.89e-10

5.33         9e-7

5.39        3.0e-10

5.56        3.3e-10

5.67        1.12e-9

5.7          5e-9

5.83       4e-10

5.92       2e-10

6.000     1.85e-7


Entry  Thu Oct 30 22:44:45 2014, Keerthi Jayamanna, Routine, General, Mass scan 3 to 4 

IOS mass scan if faulty. These is a manual scan from mass 3 to 4

Const22 = 0.04187  adjusted to mass = 4

A/q               FC6           X          Y 


3.0075        8.14e-8      398  237

3.1222        4.93e-10   398  237

3.2064        6.24e-7     398   237

3.2613        1.05e-9     408   237

3.3395        2.29e-10   408  237

3.4072        5.6e-10     408  237

3.5089        3.36e-8     408  237

3.5600        1.4e-10     408  237

3.6069        1.86-9       408  237

3.6436        8.46e-10  408  237

3.7514        4.02e-10  418   237

3.8387        4.37e-9    418   237

3.8883        3.43e-10  418   237

3.9416        1e-10       418    237

4                   1.1e-6     418    237       500  

ELOG V2.9.2-2455